# Matryoshka Quantization

### v1

@@ Appendix A Addition Training Details
Appendix B Detailed Downstream Evaluations for OmniQuant ad QAT
Tables 7, 8, 9, 10, 11, and 12 present downstream evaluation results on Gemma-2 2B, Gemma-2 9B and Mistral 7B with OmniQuant and QAT.
Appendix C Detailed Downstream Evaluations for ${\rm MatQuant}$ Re-weighting
Tables 13, 14, and 15 present downstream evaluation results for OmniQuant reweighting experiments on Gemma-2 2B, Gemma-2 9B and Mistral 7B.
Appendix D Detailed Downstream Evaluations for Co-Distillation
Tables 16 and 17 present the downstream evaluation and perplexity results and for ${\rm MatQuant}$ co-distillation on Gemma-2 9B with OmniQuant and QAT.
Appendix E Detailed Evaluations for FFN + Attention Quantization
Tables 18 and 19 present the downstream evaluation and perplexity results for FFN + Attention quantization on Gemma-2 9B and Mistral 7B with OmniQuant and QAT.
Appendix F Detailed Evaluation for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
Tables 20, 21, 22, and 23 present the downstream evaluation results comparing Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ to ${\rm MatQuant}$ and the Baseline for int2 quantization of Gemma-2 2B, Gemma-2 9B and Mistral 7B with OmniQuant and QAT. Since Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ slices 2 bits from an 8-bit model and computes loss only over the first two bits, we can evaluate the Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ model trained for 2-bits on int4 and int8. Downstream evaluation and perplexity results for this are presented in Tables 21 and 22. We also plot the weight distribution for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ in Figure 3.
@@ Appendix F Detailed Evaluation for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
int8 $(1,1,1)$ $48.04$ $71.8$ $75.78$ $67.64$ $78.07$ $63.22$ $67.42$
$(1\sqrt{2},\sqrt{2})$ $47.35$ $71.34$ $75.66$ $67.99$ $78.07$ $63.38$ $67.3$
$(\sqrt{2},1,\sqrt{2})$ $47.44$ $72.43$ $76.02$ $67.45$ $78.02$ $63.85$ $67.54$
$(1,1\sqrt{2})$ $47.7$ $71.89$ $75.63$ $67.21$ $78.07$ $63.38$ $67.31$
$(2,2,1)$ $48.38$ $72.31$ $76.3$ $68.32$ $78.35$ $63.46$ $67.85$
$(\sqrt{2},2,1)$ $48.46$ $71.84$ $75.93$ $68.35$ $77.91$ $63.14$ $67.6$
$(2,\sqrt{2},1)$ $47.95$ $71.72$ $75.26$ $68.13$ $78.07$ $62.75$ $67.31$
$(\sqrt{2},\sqrt{2},1)$ $47.35$ $71.34$ $75.66$ $67.99$ $78.07$ $63.38$ $67.3$
int4 $(1,1,1)$ $45.65$ $70.29$ $74.8$ $66.07$ $77.58$ $62.27$ $66.11$
$(1\sqrt{2},\sqrt{2})$ $46.33$ $70.92$ $73.7$ $67.67$ $77.26$ $62.9$ $66.46$
$(\sqrt{2},1,\sqrt{2})$ $46.42$ $70.96$ $74.71$ $65.78$ $77.58$ $63.14$ $66.43$
$(1,1\sqrt{2})$ $45.56$ $71.55$ $75.75$ $66.18$ $77.48$ $63.69$ $66.7$
$(2,2,1)$ $46.84$ $70.88$ $74.92$ $66.48$ $77.91$ $62.19$ $66.54$
$(\sqrt{2},2,1)$ $47.35$ $71.68$ $72.69$ $66.79$ $77.26$ $63.38$ $66.52$
$(2,\sqrt{2},1)$ $45.9$ $70.83$ $75.11$ $66.97$ $77.37$ $62.27$ $66.41$
$(\sqrt{2},\sqrt{2},1)$ $46.33$ $70.92$ $73.7$ $67.67$ $77.26$ $62.9$ $66.46$
int2 $(1,1,1)$ $34.39$ $59.64$ $62.69$ $52.11$ $69.86$ $55.56$ $55.71$
$(1\sqrt{2},\sqrt{2})$ $32.76$ $56.99$ $63.46$ $51.99$ $70.29$ $56.27$ $55.29$
$(\sqrt{2},1,\sqrt{2})$ $35.07$ $62.04$ $65.78$ $54.26$ $71.65$ $56.27$ $57.51$
$(1,1\sqrt{2})$ $34.22$ $60.4$ $64.98$ $54.3$ $71.38$ $57.22$ $57.08$
$(2,2,1)$ $34.47$ $57.95$ $63.94$ $51.84$ $69.75$ $56.27$ $55.7$
$(\sqrt{2},2,1)$ $33.45$ $57.49$ $65.02$ $52.22$ $70.4$ $55.64$ $55.7$
$(2,\sqrt{2},1)$ $34.04$ $58.84$ $65.11$ $51.77$ $70.89$ $57.14$ $56.3$
$(\sqrt{2},\sqrt{2},1)$ $32.76$ $56.99$ $63.46$ $51.99$ $70.29$ $56.27$ $55.29$
int6 $(1,1,1)$ $47.1$ $71.46$ $76.02$ $67.47$ $77.91$ $63.61$ $67.26$
$(1\sqrt{2},\sqrt{2})$ $47.44$ $71.42$ $74.95$ $67.85$ $77.86$ $63.3$ $67.14$
$(\sqrt{2},1,\sqrt{2})$ $47.61$ $71.89$ $75.9$ $67.37$ $78.24$ $63.77$ $67.46$
$(1,1\sqrt{2})$ $47.78$ $71.63$ $75.47$ $67.2$ $77.86$ $63.61$ $67.26$
$(2,2,1)$ $48.55$ $72.69$ $76.3$ $68.02$ $78.67$ $63.85$ $68.01$
$(\sqrt{2},2,1)$ $48.29$ $71.76$ $75.72$ $68.42$ $78.02$ $63.38$ $67.6$
$(2,\sqrt{2},1)$ $48.38$ $71.51$ $75.84$ $68.24$ $78.18$ $63.85$ $67.67$
$(\sqrt{2},\sqrt{2},1)$ $47.44$ $71.42$ $74.95$ $67.85$ $77.86$ $63.3$ $67.14$
int3 $(1,1,1)$ $44.45$ $68.56$ $69.11$ $62.28$ $75.95$ $62.59$ $63.82$
$(1\sqrt{2},\sqrt{2})$ $43.17$ $68.73$ $64.74$ $61.31$ $76.39$ $61.48$ $62.64$
$(\sqrt{2},1,\sqrt{2})$ $41.98$ $68.6$ $70.34$ $61.95$ $75.9$ $63.3$ $63.68$
$(1,1\sqrt{2})$ $41.64$ $66.71$ $71.62$ $61.94$ $76.01$ $61.09$ $63.17$
$(2,2,1)$ $41.98$ $68.35$ $68.41$ $63.74$ $76.17$ $60.77$ $63.24$
$(\sqrt{2},2,1)$ $42.66$ $66.54$ $70.46$ $63.61$ $75.63$ $62.98$ $63.65$
$(2,\sqrt{2},1)$ $43.17$ $66.71$ $60.03$ $62.71$ $76.77$ $61.64$ $61.84$
$(\sqrt{2},\sqrt{2},1)$ $43.17$ $68.73$ $64.74$ $61.31$ $76.39$ $61.48$ $62.64$
@@
int8 $(1,1,1)$ $58.11$ $78.03$ $83.27$ $76.17$ $81.18$ $67.09$ $73.97$
$(1\sqrt{2},\sqrt{2})$ $57.68$ $77.4$ $83.73$ $76.1$ $81.18$ $67.64$ $73.95$
$(\sqrt{2},1,\sqrt{2})$ $58.11$ $77.86$ $81.04$ $76$ $81.18$ $67.09$ $73.55$
$(1,1\sqrt{2})$ $56.91$ $77.1$ $82.39$ $75.93$ $81.18$ $67.17$ $73.45$
$(2,2,1)$ $58.79$ $77.48$ $82.66$ $76.55$ $81.23$ $67.4$ $74.02$
$(\sqrt{2},2,1)$ $58.53$ $77.31$ $82.63$ $76.54$ $80.96$ $67.56$ $73.92$
$(2,\sqrt{2},1)$ $58.62$ $77.27$ $84.31$ $76.54$ $81.34$ $66.85$ $74.16$
$(\sqrt{2},\sqrt{2},1)$ $59.13$ $78.07$ $84.16$ $76.46$ $80.9$ $67.25$ $74.33$
int4 $(1,1,1)$ $57.25$ $77.36$ $84.86$ $75.52$ $81.5$ $66.77$ $73.88$
$(1\sqrt{2},\sqrt{2})$ $56.74$ $77.74$ $85.08$ $75.5$ $80.85$ $66.85$ $73.79$
$(\sqrt{2},1,\sqrt{2})$ $57.42$ $78.28$ $82.51$ $75.97$ $81.34$ $67.56$ $73.85$
$(1,1\sqrt{2})$ $57.59$ $77.82$ $84.28$ $75.32$ $81.12$ $66.38$ $73.75$
$(2,2,1)$ $58.62$ $78.28$ $83.67$ $76.01$ $81.5$ $67.88$ $74.33$
$(\sqrt{2},2,1)$ $58.19$ $77.82$ $83.91$ $76.62$ $81.99$ $67.72$ $74.37$
$(2,\sqrt{2},1)$ $58.28$ $78.16$ $84.53$ $76.41$ $81.72$ $67.09$ $74.36$
$(\sqrt{2},\sqrt{2},1)$ $57.94$ $78.11$ $84.98$ $76.5$ $81.01$ $67.01$ $74.26$
int2 $(1,1,1)$ $48.72$ $72.18$ $79.2$ $68.11$ $76.17$ $66.77$ $68.52$
$(1\sqrt{2},\sqrt{2})$ $49.83$ $73.91$ $78.75$ $67.27$ $77.2$ $66.46$ $68.9$
$(\sqrt{2},1,\sqrt{2})$ $48.55$ $74.24$ $81.5$ $68.44$ $76.5$ $65.9$ $69.19$
$(1,1\sqrt{2})$ $48.29$ $72.94$ $74.74$ $68.34$ $77.58$ $65.67$ $67.93$
$(2,2,1)$ $46.76$ $73.27$ $71.96$ $67.98$ $76.77$ $63.61$ $66.72$
$(\sqrt{2},2,1)$ $46.76$ $73.7$ $77.65$ $67.01$ $77.58$ $65.98$ $68.11$
$(2,\sqrt{2},1)$ $46.76$ $72.35$ $75.35$ $67.51$ $76.39$ $67.56$ $67.65$
$(\sqrt{2},\sqrt{2},1)$ $46.59$ $72.6$ $79.3$ $67.58$ $77.69$ $65.75$ $68.25$
int6 $(1,1,1)$ $58.87$ $78.03$ $83.61$ $76.18$ $81.45$ $67.09$ $74.21$
$(1\sqrt{2},\sqrt{2})$ $57.51$ $77.53$ $83.55$ $75.98$ $80.9$ $67.17$ $73.77$
$(\sqrt{2},1,\sqrt{2})$ $58.79$ $77.82$ $81.38$ $76.21$ $81.07$ $67.72$ $73.83$
$(1,1\sqrt{2})$ $57.34$ $77.23$ $82.57$ $75.89$ $81.12$ $67.17$ $73.55$
$(2,2,1)$ $59.04$ $77.4$ $82.66$ $76.55$ $81.56$ $68.03$ $74.21$
$(\sqrt{2},2,1)$ $59.22$ $77.65$ $82.17$ $76.62$ $81.23$ $67.8$ $74.11$
$(2,\sqrt{2},1)$ $58.36$ $77.82$ $83.79$ $76.47$ $81.23$ $67.25$ $74.15$
$(\sqrt{2},\sqrt{2},1)$ $59.3$ $78.37$ $84.5$ $76.57$ $80.85$ $67.4$ $74.5$
int3 $(1,1,1)$ $55.46$ $76.14$ $84.04$ $74.49$ $80.14$ $67.32$ $72.93$
$(1\sqrt{2},\sqrt{2})$ $56.23$ $76.05$ $82.6$ $74.85$ $80.9$ $67.01$ $72.94$
$(\sqrt{2},1,\sqrt{2})$ $56.4$ $77.86$ $80.64$ $75.11$ $79.87$ $68.51$ $73.06$
$(1,1\sqrt{2})$ $55.63$ $76.05$ $82.39$ $74.21$ $80.3$ $67.17$ $72.62$
$(2,2,1)$ $55.2$ $76.56$ $84.19$ $74.87$ $80.2$ $67.72$ $73.12$
$(\sqrt{2},2,1)$ $54.44$ $75.63$ $80.55$ $74.97$ $80.96$ $67.72$ $72.38$
$(2,\sqrt{2},1)$ $56.14$ $75.67$ $83.33$ $74.96$ $80.52$ $67.72$ $73.06$
$(\sqrt{2},\sqrt{2},1)$ $56.31$ $77.4$ $83.24$ $75.62$ $80.41$ $66.54$ $73.25$
@@
int8 $(1,1,1)$ $48.04$ $73.44$ $84.13$ $79.37$ $81.12$ $74.66$ $73.46$
$(1\sqrt{2},\sqrt{2})$ $48.46$ $73.19$ $84.28$ $79.19$ $81.12$ $74.74$ $73.5$
$(\sqrt{2},1,\sqrt{2})$ $47.95$ $73.4$ $84.46$ $79.11$ $81.34$ $74.51$ $73.46$
$(1,1\sqrt{2})$ $48.21$ $73.02$ $84.34$ $79.03$ $81.28$ $74.59$ $73.41$
$(2,2,1)$ $49.06$ $73.48$ $84.74$ $79.73$ $81.56$ $74.35$ $73.82$
$(\sqrt{2},2,1)$ $49.06$ $73.57$ $84.56$ $79.64$ $81.39$ $74.27$ $73.75$
$(2,\sqrt{2},1)$ $48.98$ $73.95$ $84.50$ $79.60$ $81.61$ $74.90$ $73.92$
$(\sqrt{2},\sqrt{2},1)$ $48.98$ $73.86$ $84.56$ $79.55$ $81.23$ $74.74$ $73.82$
int4 $(1,1,1)$ $48.21$ $72.69$ $83.49$ $78.82$ $81.12$ $74.43$ $73.13$
$(1\sqrt{2},\sqrt{2})$ $49.15$ $72.81$ $83.39$ $78.71$ $80.79$ $74.66$ $73.25$
$(\sqrt{2},1,\sqrt{2})$ $47.95$ $72.43$ $83.43$ $79.24$ $81.01$ $74.03$ $73.01$
$(1,1\sqrt{2})$ $48.46$ $73.44$ $84.07$ $78.9$ $81.01$ $73.88$ $73.29$
$(2,2,1)$ $49.15$ $72.81$ $83.88$ $79.8$ $81.88$ $73.48$ $73.5$
$(\sqrt{2},2,1)$ $48.89$ $72.69$ $82.72$ $79.53$ $81.66$ $73.88$ $73.23$
$(2,\sqrt{2},1)$ $47.87$ $72.05$ $83$ $79.56$ $81.23$ $74.27$ $73$
$(\sqrt{2},\sqrt{2},1)$ $48.29$ $72.47$ $82.84$ $79.52$ $81.07$ $73.64$ $72.97$
int2 $(1,1,1)$ $41.38$ $67.42$ $71.62$ $71.98$ $77.86$ $65.67$ $65.99$
$(1\sqrt{2},\sqrt{2})$ $40.78$ $66.2$ $73.61$ $72.68$ $77.75$ $67.4$ $66.4$
$(\sqrt{2},1,\sqrt{2})$ $40.36$ $67.09$ $75.35$ $72.46$ $77.48$ $65.9$ $66.44$
$(1,1\sqrt{2})$ $40.36$ $67.17$ $74.83$ $71.64$ $77.53$ $66.14$ $66.28$
$(2,2,1)$ $37.2$ $62.46$ $67.74$ $70.29$ $76.55$ $66.69$ $63.49$
$(\sqrt{2},2,1)$ $37.29$ $64.35$ $61.1$ $68.88$ $74.86$ $65.19$ $61.94$
$(2,\sqrt{2},1)$ $39.68$ $65.24$ $68.93$ $66.64$ $75.19$ $64.09$ $63.29$
$(\sqrt{2},\sqrt{2},1)$ $34.56$ $61.24$ $60.61$ $58.07$ $72.63$ $59.98$ $57.85$
int6 $(1,1,1)$ $48.46$ $72.98$ $84.07$ $79.64$ $81.18$ $75.22$ $73.59$
$(1\sqrt{2},\sqrt{2})$ $49.06$ $73.44$ $84.59$ $79.51$ $81.28$ $74.74$ $73.77$
$(\sqrt{2},1,\sqrt{2})$ $47.95$ $73.48$ $84.43$ $79.28$ $81.45$ $75.14$ $73.62$
$(1,1\sqrt{2})$ $48.38$ $72.94$ $84.34$ $79.15$ $81.18$ $74.59$ $73.43$
$(2,2,1)$ $48.46$ $72.94$ $84.13$ $79.89$ $81.5$ $74.9$ $73.64$
$(\sqrt{2},2,1)$ $48.81$ $73.48$ $84.34$ $79.67$ $81.34$ $74.9$ $73.76$
$(2,\sqrt{2},1)$ $49.4$ $73.65$ $84.4$ $79.68$ $81.28$ $74.74$ $73.86$
$(\sqrt{2},\sqrt{2},1)$ $49.23$ $73.57$ $84.43$ $79.55$ $81.12$ $74.66$ $73.76$
int3 $(1,1,1)$ $45.65$ $71.21$ $80.43$ $78.31$ $81.07$ $72.61$ $71.55$
$(1\sqrt{2},\sqrt{2})$ $47.7$ $72.05$ $82.81$ $78.74$ $81.12$ $72.77$ $72.53$
$(\sqrt{2},1,\sqrt{2})$ $46.33$ $72.43$ $81.8$ $79.03$ $82.1$ $73.4$ $72.51$
$(1,1\sqrt{2})$ $45.99$ $71.09$ $80.73$ $78.77$ $80.85$ $72.53$ $71.66$
$(2,2,1)$ $47.95$ $73.36$ $82.57$ $79.31$ $81.39$ $74.9$ $73.25$
$(\sqrt{2},2,1)$ $44.45$ $69.7$ $82.11$ $77.68$ $80.2$ $71.74$ $70.98$
$(2,\sqrt{2},1)$ $46.84$ $72.73$ $80.95$ $78.79$ $81.56$ $73.01$ $72.31$
$(\sqrt{2},\sqrt{2},1)$ $47.01$ $71.59$ $81.96$ $78.89$ $81.39$ $72.45$ $72.22$

### v2

@@ Appendix B Addition Training Details
Appendix C Detailed Downstream Evaluations for OmniQuant and QAT
Tables 9, 10, 11, 12, 13, and 14 present downstream evaluation results on Gemma-2 2B, Gemma-2 9B and Mistral 7B with OmniQuant and QAT.
Appendix D Detailed Downstream Evaluations for ${\rm MatQuant}$ Re-weighting
Tables 15, 17, and 16 present downstream evaluation results for OmniQuant reweighting experiments on Gemma-2 2B, Gemma-2 9B and Mistral 7B.
Appendix E Detailed Downstream Evaluations for Co-Distillation
Tables 18 and 19 present the downstream evaluation and perplexity results for ${\rm MatQuant}$ with co-distillation on Gemma-2 9B. We present results with both, OmniQuant and QAT as the base algorithms.
Appendix F Detailed Evaluations for FFN + Attention Quantization
Tables 20 and 21 present the downstream evaluation and perplexity results for FFN + Attention quantization on Gemma-2 9B and Mistral 7B with OmniQuant and QAT.
Appendix G Detailed Evaluation for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
Tables 22, 23, 24, and 25 present the downstream evaluation results comparing Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ to ${\rm MatQuant}$ and the Baseline for int2 quantization of Gemma-2 2B, Gemma-2 9B and Mistral 7B with OmniQuant and QAT. Since Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ slices 2 bits from an 8-bit model and computes loss only over the first two bits, we can evaluate the Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ model trained for 2-bits on int4 and int8. Downstream evaluation and perplexity results for this are presented in Tables 23 and 24. We also plot the weight distribution for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ in Figure 4.
@@ Appendix G Detailed Evaluation for Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
Appendix H Detailed Evaluation for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$
Tables 26, 27, and 28 present downstream evaluation results for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ when applied to Gemma-2 2B, 9B, and Mistral 7B with OmniQuant as the base algorithm. Table 29 presents downstream evaluation and perplexity results for our Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ co-distillation experiments on Gemma-2 9B with OmniQuant as the base algorithm.
@@ Appendix H Detailed Evaluation for Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$
int8 $(0.1,0.1,1)$ $49.66$ $71$ $76.73$ $68.85$ $78.56$ $63.3$ $68.02$
$(0.2,0.2,1)$ $49.4$ $71.3$ $76.21$ $68.97$ $78.29$ $63.3$ $67.91$
$(0.3,0.3,1)$ $48.81$ $71.72$ $76.57$ $68.95$ $78.4$ $63.61$ $68.01$
$(0.4,0.4,1)$ $48.72$ $71.72$ $76.61$ $68.92$ $78.73$ $62.98$ $67.95$
$(0.5,0.5,1)$ $49.06$ $71.34$ $76.15$ $68.86$ $78.45$ $62.98$ $67.81$
int4 $(0.1,0.1,1)$ $47.27$ $70.79$ $73.76$ $66.85$ $78.07$ $62.75$ $66.58$
$(0.2,0.2,1)$ $48.63$ $71$ $76.06$ $68.11$ $77.97$ $63.06$ $67.47$
$(0.3,0.3,1)$ $47.7$ $71.17$ $75.08$ $67.57$ $77.69$ $62.59$ $66.97$
$(0.4,0.4,1)$ $48.29$ $71.25$ $76.76$ $67.46$ $77.58$ $63.54$ $67.48$
$(0.5,0.5,1)$ $48.04$ $70.66$ $75.9$ $67.57$ $78.4$ $64.01$ $67.43$
int2 $(0.1,0.1,1)$ $29.95$ $54.21$ $64.4$ $44.37$ $66.81$ $54.46$ $52.37$
$(0.2,0.2,1)$ $30.03$ $52.78$ $62.39$ $44.66$ $66.81$ $54.62$ $51.88$
$(0.3,0.3,1)$ $29.18$ $52.61$ $62.57$ $41.41$ $65.94$ $54.62$ $51.05$
$(0.4,0.4,1)$ $28.75$ $54.88$ $62.17$ $42.53$ $66.16$ $55.64$ $51.69$
$(0.5,0.5,1)$ $27.13$ $51.05$ $60.95$ $39.94$ $65.56$ $54.3$ $49.82$
int6 $(0.1,0.1,1)$ $48.89$ $70.5$ $75.69$ $68.89$ $78.4$ $62.75$ $67.52$
$(0.2,0.2,1)$ $49.32$ $70.96$ $75.87$ $68.93$ $78.29$ $62.67$ $67.67$
$(0.3,0.3,1)$ $48.98$ $71.63$ $76.21$ $68.68$ $78.73$ $63.46$ $67.95$
$(0.4,0.4,1)$ $48.98$ $71.72$ $75.75$ $68.83$ $78.67$ $63.61$ $67.93$
$(0.5,0.5,1)$ $49.4$ $71.59$ $76.21$ $68.63$ $78.29$ $63.85$ $67.99$
int3 $(0.1,0.1,1)$ $44.03$ $67.09$ $74.25$ $62.78$ $77.26$ $61.4$ $64.47$
$(0.2,0.2,1)$ $43.09$ $65.7$ $67.19$ $59.57$ $75.3$ $60.38$ $61.87$
$(0.3,0.3,1)$ $43.94$ $68.35$ $71.87$ $59.54$ $75.79$ $59.98$ $63.24$
$(0.4,0.4,1)$ $41.81$ $65.53$ $72.91$ $61.42$ $75.03$ $61.88$ $63.1$
$(0.5,0.5,1)$ $41.64$ $67.34$ $71.87$ $61.15$ $74.54$ $61.64$ $63.03$
@@
int8 $(0.1,0.1,1)$ $57.59$ $77.02$ $84.01$ $76.61$ $81.18$ $67.88$ $74.05$
$(0.2,0.2,1)$ $57.76$ $76.73$ $83.73$ $76.5$ $81.34$ $67.4$ $73.91$
$(0.3,0.3,1)$ $57.94$ $76.64$ $83.36$ $76.56$ $81.01$ $67.8$ $73.88$
$(0.4,0.4,1)$ $58.28$ $76.52$ $83.15$ $76.74$ $80.96$ $67.4$ $73.84$
$(0.5,0.5,1)$ $57.68$ $76.68$ $83.39$ $76.62$ $81.07$ $67.09$ $73.75$
int4 $(0.1,0.1,1)$ $58.02$ $78.11$ $83.24$ $76.08$ $80.96$ $66.54$ $73.83$
$(0.2,0.2,1)$ $58.96$ $77.9$ $82.57$ $76.14$ $81.07$ $66.14$ $73.8$
$(0.3,0.3,1)$ $57.42$ $77.23$ $81.62$ $75.72$ $80.85$ $66.69$ $73.25$
$(0.4,0.4,1)$ $58.96$ $78.32$ $84.53$ $76.17$ $81.45$ $66.46$ $74.32$
$(0.5,0.5,1)$ $57.08$ $77.02$ $84.65$ $76.11$ $81.56$ $66.06$ $73.75$
int2 $(0.1,0.1,1)$ $40.78$ $67.85$ $73.64$ $60.56$ $72.09$ $65.19$ $63.35$
$(0.2,0.2,1)$ $40.53$ $67.97$ $75.57$ $60.83$ $72.25$ $67.09$ $64.04$
$(0.3,0.3,1)$ $39.42$ $67.68$ $79.08$ $60.79$ $72.47$ $65.19$ $64.1$
$(0.4,0.4,1)$ $39.68$ $66.54$ $66.24$ $61.08$ $73.07$ $65.27$ $61.98$
$(0.5,0.5,1)$ $40.02$ $66.16$ $69.08$ $60.54$ $73.23$ $64.88$ $62.32$
int6 $(0.1,0.1,1)$ $57.25$ $76.94$ $84.04$ $76.63$ $81.34$ $67.32$ $73.92$
$(0.2,0.2,1)$ $57.25$ $76.6$ $83.79$ $76.46$ $81.12$ $67.64$ $73.81$
$(0.3,0.3,1)$ $58.7$ $76.98$ $83.09$ $76.63$ $80.69$ $67.32$ $73.9$
$(0.4,0.4,1)$ $58.28$ $76.43$ $83.15$ $76.76$ $81.18$ $67.09$ $73.81$
$(0.5,0.5,1)$ $58.28$ $76.3$ $83.33$ $76.68$ $81.18$ $66.93$ $73.78$
int3 $(0.1,0.1,1)$ $55.8$ $76.89$ $81.99$ $74.27$ $80.14$ $68.11$ $72.87$
$(0.2,0.2,1)$ $54.69$ $76.56$ $79.79$ $73.92$ $79.92$ $66.77$ $71.94$
$(0.3,0.3,1)$ $56.48$ $77.53$ $83.09$ $73.71$ $80.69$ $67.32$ $73.14$
$(0.4,0.4,1)$ $56.23$ $77.86$ $83.79$ $74.12$ $80.69$ $68.98$ $73.61$
$(0.5,0.5,1)$ $54.35$ $76.3$ $83.67$ $74.21$ $80.09$ $68.03$ $72.77$
@@
int8 $(0.1,0.1,1)$ $49.23$ $71.84$ $83.94$ $78.9$ $81.39$ $74.35$ $73.27$
$(0.2,0.2,1)$ $49.23$ $71.97$ $83.91$ $79.04$ $81.5$ $74.98$ $73.44$
$(0.3,0.3,1)$ $49.32$ $72.39$ $84.43$ $79.24$ $81.23$ $74.74$ $73.56$
$(0.4,0.4,1)$ $49.06$ $72.52$ $84.74$ $79.21$ $81.45$ $74.9$ $73.65$
$(0.5,0.5,1)$ $49.15$ $72.64$ $84.65$ $79.37$ $81.72$ $74.82$ $73.72$
int4 $(0.1,0.1,1)$ $47.61$ $71.59$ $83.3$ $78.32$ $81.61$ $74.11$ $72.76$
$(0.2,0.2,1)$ $48.12$ $72.14$ $84.07$ $78.72$ $81.45$ $74.43$ $73.16$
$(0.3,0.3,1)$ $48.21$ $72.81$ $84.4$ $79.02$ $81.18$ $75.22$ $73.47$
$(0.4,0.4,1)$ $47.87$ $71.55$ $83.88$ $78.85$ $81.34$ $74.9$ $73.06$
$(0.5,0.5,1)$ $48.21$ $71.97$ $83.82$ $79.03$ $81.39$ $74.35$ $73.13$
int2 $(0.1,0.1,1)$ $37.46$ $63.43$ $71.53$ $66.22$ $75.24$ $65.59$ $63.25$
$(0.2,0.2,1)$ $37.54$ $64.81$ $71.8$ $66.57$ $74.37$ $65.27$ $63.39$
$(0.3,0.3,1)$ $37.46$ $62.92$ $75.35$ $67.2$ $74.43$ $64.25$ $63.6$
$(0.4,0.4,1)$ $37.88$ $62.58$ $73.15$ $65.89$ $73.88$ $63.14$ $62.75$
$(0.5,0.5,1)$ $37.29$ $62.75$ $69.36$ $64.99$ $72.36$ $64.25$ $61.83$
int6 $(0.1,0.1,1)$ $49.57$ $71.72$ $83.76$ $78.87$ $81.28$ $74.03$ $73.2$
$(0.2,0.2,1)$ $49.49$ $72.52$ $84.22$ $79.08$ $81.39$ $74.19$ $73.48$
$(0.3,0.3,1)$ $48.89$ $72.01$ $83.85$ $79.2$ $81.39$ $74.35$ $73.28$
$(0.4,0.4,1)$ $49.4$ $72.47$ $84.68$ $79.52$ $81.34$ $74.35$ $73.63$
$(0.5,0.5,1)$ $49.4$ $72.39$ $84.31$ $79.5$ $81.28$ $74.27$ $73.52$
int3 $(0.1,0.1,1)$ $44.88$ $68.22$ $81.96$ $76.13$ $80.69$ $71.35$ $70.54$
$(0.2,0.2,1)$ $43.94$ $67.85$ $81.56$ $76.55$ $79.76$ $72.61$ $70.38$
$(0.3,0.3,1)$ $45.39$ $67.89$ $80.92$ $77.13$ $80.47$ $72.06$ $70.64$
$(0.4,0.4,1)$ $47.35$ $71$ $80$ $76.96$ $80.3$ $71.35$ $71.16$
$(0.5,0.5,1)$ $46.76$ $70.29$ $82.17$ $77.32$ $80.9$ $71.11$ $71.43$